According to a newspaper account, a paratrooper survived a training jump from 1214 ft when his parachute failed to open but provided some resistance by flapping in the wind. Allegedly he hit the ground at 97.5 mih after falling for 7 seconds. To test the accuracy of this account, you should first find the drag coefficient p, assuming a terminal velocity of 97.5 mi/h and also that the resistance of the paratrooper falling through the air is proportional to his velocity. Remember that the accleration due to gravity near the earths surface is 32 ft/sec2 sec-l Next, find the distance D fallen in 7 seconds is the newspaper account reasonable? Enter Y or N
Answer #1

We take downward direction as positive in the problem

now,

acceleration is given by   dv dt

according to the problem

rac{dv}{dt} = g - ho v

or

dt

The integrating factor is opt therefore

(eptu) = get dt

ot et C

or

(t) -(1 -ept) eq-1

we know that the terminal velocity is 97.5 m/h or 143 ft/s

as time tends to infinity  e^{ ho t} = 0  

by plugging all the values we get

143ーシ(1-0)

or

32 ρ-143 (as g=32 ft/s2 is given.)

thus we get

ρ-0.2238s-1

Part-b

Again using equation -1 in y direction displacement over time we get

ty = u(t) = g(1-et)

ot dt

or

y(t)= (rac{g}{ ho})(t+ (rac{1}{ ho})e^{(- ho t)}) + K eq-2   (integrating w.r.t time)

we know that at   t=0y(0) 0

plugging these values in eq-2 we get

K = -(rac{g}{ ho ^2})

or

32 (0.2238)2 K =-( )

which gives us

Kー-638.9 ft

thus the eq-2 will become

y(t)-(0.2235 32 e-(0.2238)t)) 638.9 0,2238

For the distance travelled in 7 seconds we have to put t=7 in above equation, therefore

e(0.22387)) 638.9

we get

y(7) = 495·365ft

Hence the distance fallen in 7 seconds is

D 495.365ft